Suppose that 70% of all dialysis patients will survive for at least 5 years.

In a simple random sample of 100 new dialysis patients, what is the probability that the proportion surviving for at least five years will exceed 80%, rounded to 5 decimal places?
Answer #1

We know that the sample proportion is distributed as:

p \sim N(P,\frac{P(1-P)}{n}) \sim N(0.7,\frac{0.7(1-0.7)}{100}) \sim N(0.7,0.0021)

Required probability =

\\P(p>0.8)=P(Z>\frac{0.8-0.7}{\sqrt{0.0021}})=P(Z>2.18)=1-P(Z<2.18)\\ \\ =1-0.98537=0.01463